OpenStudy (anonymous):

solve ∠ABC for the missing values. Give: m∠B=22° , b=16.8 , a=22.42 case1: m∠A= m∠C= c= Case2: M∠A= M∠C= c=

OpenStudy (campbell_st):

do you know the law of sine..?

OpenStudy (anonymous):

is there a pic?

OpenStudy (campbell_st):

the picture doesn't match either of the question..

OpenStudy (campbell_st):

so to solve the original questions, do you know the law of sines..?

OpenStudy (anonymous):



OpenStudy (anonymous):

makes more sense?

OpenStudy (anonymous):

 does this look right so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

|just plug in

OpenStudy (anonymous):

so A/16=22/22.4 A=16.29

OpenStudy (anonymous):

does that look right @campbell_st

OpenStudy (anonymous):

what math are you in?

OpenStudy (anonymous):

theres case 1 and case 2 and is pre-calc

OpenStudy (anonymous):

oh well thats why it was hard i am in geomatry

OpenStudy (anonymous):

im prob wrong but it does have something to do with sine rule

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay